### v1

# SKETCH: Semantic Key-Point Conditioning for Long-Horizon Vessel Trajectory Prediction

Abstract
Accurate long-horizon vessel trajectory prediction remains challenging due to compounded uncertainty from complex navigation behaviors and environmental factors. Existing methods often struggle to maintain global directional consistency, leading to drifting or implausible trajectories when extrapolated over long time horizons. To address this issue, we propose a semantic-key-point-conditioned trajectory modeling framework, in which future trajectories are predicted by conditioning on a high-level Next Key Point (NKP) that captures navigational intent. This formulation decomposes long-horizon prediction into global semantic decision-making and local motion modeling, effectively restricting the support of future trajectories to semantically feasible subsets. To efficiently estimate the NKP prior from historical observations, we adopt a pretrain-finetune strategy. Extensive experiments on real-world AIS data demonstrate that the proposed method consistently outperforms state-of-the-art approaches, particularly for long travel durations, directional accuracy, and fine-grained trajectory prediction.
Navigational Intent, Next Key Point, Semantic Conditioning, Large Language Model, Long-Horizon Trajectory Prediction
1 Introduction
Maritime transportation is the basis of global trade, carrying over 80% of goods worldwide by volume (Kosowska-Stamirowska, 2020). Accurate vessel trajectory prediction is therefore a fundamental capability for maritime intelligence, supporting collision avoidance (Papadimitrakis et al., 2021; Zhang et al., 2022), port operation optimization (Peng et al., 2023), search and rescue (Liu et al., 2024), and fuel-efficient voyage planning (Zhang et al., 2025). Meanwhile, the widespread use of the Automatic Identification System (AIS) has enabled data-driven approaches by providing large-scale, high-frequency vessel motion records (Hexeberg et al., 2017; Murray and Perera, 2020; Wang et al., 2023b; Yang et al., 2022). However, the availability of large-scale data alone does not eliminate the intrinsic difficulty of long-horizon trajectory forecasting, where uncertainty compounds over time.
Building upon such data, existing approaches can be broadly categorized into non-learning and learning-based methods. Early non-learning approaches embed explicit mathematical models or statistical priors, including the Constant Velocity Model (CVM) (Xiao et al., 2020), the Nearly Constant Velocity (NCV) model (Ristic et al., 2008), Extended Kalman Filter (EKF) variants (Perera et al., 2012; Perera and Soares, 2010), stochastic processes such as the Ornstein–Uhlenbeck (OU) process (Millefiori et al., 2017), Gaussian Process (GP) models (Rong et al., 2019), and knowledge-based particle filtering methods (Mazzarella et al., 2015). These approaches are favored for their computational efficiency and simplicity (Chen et al., 2023; Xiao et al., 2020), but rely on simplified abstractions of complex vessel dynamics that often neglect environmental influences (Kanazawa et al., 2021). As a result, their performance degrades significantly during maneuvers and under noisy conditions (Chen et al., 2023; Gao et al., 2021), making them inherently inadequate for medium- and long-horizon trajectory prediction (Nguyen and Fablet, 2024).
To overcome the limitations of purely statistical modeling, learning-based approaches have been widely explored for sequential trajectory prediction. Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks improve expressive capacity but often struggle to capture long-range dependencies and multimodal future behaviors (Gao et al., 2021; Chen et al., 2023). In particular, MP-LSTM predicts midpoints and last points, and interpolates intermediate trajectories, which limits their ability to model fine-grained kinematic variations and directional changes (Gao et al., 2021). These limitations become increasingly pronounced as the prediction horizon grows.
More recently, Transformer-based models have shown promise due to their ability to model long-range dependencies via self-attention (Vaswani et al., 2017). Approaches such as TrAISformer enhance medium-term trajectory prediction and more effectively capture multimodal uncertainty (Nguyen and Fablet, 2024). Nevertheless, we observe that standard autoregressive Transformers still degrade severely in long-horizon settings. Notably, both TrAISformer and recent token-based GPT-style forecasters, such as TrackGPT (Stroh, 2024), convert continuous trajectory attributes into discrete tokens and learn embeddings over these discrete representations as the input to a Transformer backbone, which will prevent unseen embedded layer regions from being trained and will be fail when trajectories enter those regions. As uncertainty accumulates, predictions tend to collapse toward nearly horizontal trajectories. This observation suggests that increasing model capacity alone is insufficient for robust long-term trajectory forecasting.
One more limitations stem from the lack of explicit modeling of global navigational way intent. In real-world maritime navigation, vessels follow a hierarchical decision process: high-level route planning toward the Next Key Points, coupled with low-level continuous control of speed and heading (Potočnik, 2025). Most existing trajectory prediction models, however, operate solely at the local kinematic level. As a result, long-term navigational intent must be implicitly inferred from short-term motion patterns, which becomes increasingly unreliable over extended horizons.
Motivated by this observation, we reformulate long-horizon vessel trajectory prediction as a hierarchical forecasting problem with explicit modeling of global navigational intent. Our main contributions are summarized as follows:
  1. 1.
    We identify the lack of explicit modeling of global navigational intent as a fundamental limitation in existing long-horizon vessel trajectory prediction methods.
  2. 2.
    We introduce the Next Key Point (NKP) as a semantic intent variable and condition trajectory prediction on it, thereby restructuring long-horizon forecasting into a hierarchical problem that separates global navigational decisions from local motion modeling.
  3. 3.
    We develop an efficient training strategy to estimate the NKP-conditioned model, enabling generalization to open-set navigational targets without relying on a fixed, closed-set of predefined geometric waypoints commonly used in prior work.
  4. 4.
    Extensive experiments on large-scale AIS datasets demonstrate state-of-the-art performance, particularly for long-duration trajectory prediction.
2 Related Works
2.1 Decoder-only Transformer Trajectory Predictor
Decoder-only Transformers, such as GPT-2 (Radford et al., 2019; Vaswani et al., 2017), consist of multiple blocks of masked multi-head self-attention and position-wise feed-forward layers, with residual connections, layer normalization, and dropout. They generate outputs autoregressively, predicting the next token conditioned on previous tokens:
$p(x_{1:T})=\prod_{t=1}^{T}p(x_{t}\mid x_{<t}).$ (1)
Originally designed for natural language, decoder-only Transformers have also been successfully applied to sequential data in computer vision (Ramesh et al., 2021), audio generation (Wang et al., 2023a), and time series forecasting (Ansari et al., 2024). In our work, vessel trajectories are treated as sequences of tokens (latitude, longitude, SOG, COG)($\text${latitude, longitude, SOG, COG}), allowing autoregressive modeling of future positions in a vessel trajectory context.
2.2 Retrieval-Augmented Verification
Retrieval-Augmented Detection (Kang et al., 2024) and Verification (Wang et al., 2025) extend traditional prediction by incorporating reference samples from an external database, effectively transforming prediction into a verification task. A query sample is compared with stored embeddings via similarity metrics, and labels are inferred from retrieved matches. This framework naturally supports open-set recognition and interpretable predictions.
We adopt this framework for NKP prediction: trajectories that are semantically similar (i.e., leading to the same Next Key Point) are represented as hidden states, and label prediction is achieved by verifying similarity with stored trajectories:
$\{\hat{h}_{k}\}_{k=1}^{K}=\arg\max_{h\in H}\text{sim}_{\cos}(h_{k},h)=\arg\max_{h\in H}\frac{h_{k}\cdot h}{\|h_{k}\|\cdot\|h\|}$ (2)
The top-$K$ retrieved trajectories guide label assignment, providing robustness against noisy or missing destination information.
2.3 Contrastive Learning for Embedding Consistency
Contrastive learning (Hadsell et al., 2006) is used to ensure that trajectories leading to the same NKP have similar embeddings while different ones are pushed apart. For hidden states $h_{1},h_{2}\in\mathbb{R}^{B\times T\times H}$, average pooling and L2 normalization yield sequence-level embeddings $\hat{v}_{1},\hat{v}_{2}$, and the cosine similarity $D_{i}=\text{sim}_{\cos}(\hat{v}_{1},\hat{v}_{2})$ is computed. The loss function is:
$\mathcal{L}_{\text{TCL}}=\frac{1}{B}\sum_{i=1}^{B}\big[(1-y_{i})D_{i}^{2}+y_{i}\max(0,M-D_{i})^{2}\big],$ (3)
where $y_{i}=1$ if the pair shares the same NKP label, otherwise $y_{i}=0$, and $M$ is a margin.
Figure 1: Overall Architecture. The inputs of AIS data go through Encoder 1 and MiniMind model 1 to be transformed into hidden state 1. The hidden state 1 will be sent into an MLP to predict the Next Key Point information. These coordinates will be sent to the encoder 2 to derive the hidden state 2. Then, hidden states 1 and 2 will be concatenated and passed through a dense layer, MiniMind model 2, and a decoder to obtain the predicted SOG and COG.
3 Methodology
3.1 Overview
To model vessel trajectories, we formulate the problem probabilistically. Let $X$ denote the observed historical trajectory, $Z$ the Next Key Point (NKP), and $Y$ the future trajectory. Our goal is to model the conditional distribution $p(Y\mid X)$, which we factorize as
$p(Y\mid X)=\sum_{Z}p(Y\mid X,Z)\,p(Z\mid X),$ (4)
where $p(Z\mid X)$ captures the uncertainty of the Next Key Point, and $p(Y\mid X,Z)$ models the future trajectory given a specific NKP.
Conditioning on $Z$ refines future trajectory prediction by introducing high-level semantic constraints that restrict the support of admissible trajectories, i.e.,
$\mathrm{supp}\big(p(Y\mid X,Z=z)\big)\subsetneq\mathrm{supp}\big(p(Y\mid X)\big).$ (5)
This refinement should be understood as a reduction of predictive uncertainty rather than a pointwise increase of posterior likelihoods (see Appendix A).
NKP represents a semantic navigational constraint that defines a subset of feasible future trajectories rather than a specific spatial anchor. This factorization enables the model to disentangle high-level navigation decisions from low-level dynamics.
In addition, we formulate NKP inference in an open-set manner, allowing the model to generalize to newly emerging navigational targets. The concrete inference mechanism is detailed in Section 3.2, and the overall architecture is illustrated in Fig. 1.
3.2 Semantic Next-Key-Point Modeling
Definition 3.1 (Semantic Navigational Key Point (NKP)).
A Semantic Navigational Key Point (NKP) is defined as a latent variable that represents high-level navigational intent of a vessel, rather than a deterministic geometric waypoint. Specifically, an NKP corresponds to an equivalence class of future trajectories that share the same navigational semantics, such as passing through a specific channel, approaching a port region, or following a major shipping lane.
To incorporate Next Key Point (NKP) information, we explicitly estimate the conditional distribution $p(Z\mid X)$, where $X$ denotes the observed trajectory history and $Z$ the corresponding NKP. Direct NKP classification scales poorly with dynamic maritime landscapes and cannot handle unseen key points, motivating a retrieval-based verification formulation. We adopt a contrastive verification-based formulation inspired by retrieval-augmented methods (Kang et al., 2024), where destination inference is augmented by retrieving and verifying semantically similar trajectories from a reference database, which is well-suited for modeling semantic consistency among trajectories leading to the same destination.
As illustrated in Fig. 2, we construct training pairs consisting of trajectories associated with either the same or different NKPs. Each trajectory is encoded into a latent representation, and the cosine similarity between two representations is used to assess whether they correspond to the same NKP. Following the contrastive learning paradigm (Hadsell et al., 2006), we optimize the loss in Eq. (2), which encourages trajectories leading to the same NKP to share similar representations while separating those associated with different NKPs.
To improve efficiency, we freeze Encoder 1 and MiniMind Model 1 after pretraining and only fine-tune a lightweight multi-layer perceptron (MLP) for NKP verification. In this way, the learned representation acts as a compact semantic embedding, capturing high-level navigational intent. In practice, we approximate marginalization over $Z$ using a MAP estimate, which performs robustly in our experiments.
During inference, an unknown trajectory is compared against a database of reference trajectories. For each reference trajectory whose similarity exceeds a predefined threshold, a vote is cast for its associated NKP. The NKP with the highest vote count is selected as the prediction, as summarized in Algorithm 1. This voting procedure can be interpreted as a non-parametric approximation of $p(Z\mid X)$ through nearest-neighbor evidence aggregation. The overall training and inference pipeline is shown in Fig. 2.
Algorithm 1 Next Key Point (NKP) Prediction
0: Observed trajectory $X$
0: Reference database $\mathcal{D}=\{(X_{i},z_{i})\}_{i=1}^{N}$
0: Similarity threshold $\tau$
0: Predicted Next Key Point $\hat{z}$
1: Compute embedding $h=f(X)$
2: Initialize vote counts $c(z)\leftarrow 0,\ \forall z$
3:for $(X_{i},z_{i})\in\mathcal{D}$ do
4:  Compute embedding $h_{i}=f(X_{i})$
5:  if $\mathrm{sim}(h,h_{i})\geq\tau$ then
6:   $c(z_{i})\leftarrow c(z_{i})+1$
7:  end if
8:end for
[…]
10:return $\hat{z}$
Figure 2: Key-Point Prediction Training Paradigm. Contrastive Learning is used to derive the hidden states of each trajectory, thereby decoupling the NKP information. To be more efficient, the two blocks trained previously are frozen and reused for fine-tuning.
3.3 Multi-stage Learning for NKP-Trajectory Modeling
This learning procedure provides a structured approximation to the joint distribution $p(Y,Z\mid X)$. As illustrated in Fig. 1, we adopt a three-stage learning framework that provides a structured approximation to the joint distribution $p(Y,Z\mid X)$, where $X$ denotes the observed trajectory history, $Z$ the Next Key Point (NKP), and $Y$ the future trajectory. The first two stages focus on learning the conditional components of the model, while the final stage performs integrated inference.
3.3.1 Conditional Trajectory Modeling (blue)
In the first stage, we train the trajectory prediction module to model the conditional distribution $p(Y\mid X,Z)$, where the ground-truth NKP is provided as oracle conditioning. This design decouples high-level navigational intent from low-level vessel dynamics, allowing the model to focus on long-horizon motion patterns under a fixed semantic constraint.
Although the Next Key Point $Z$ is unknown at inference time, we first learn $p(Y\mid X,Z)$ under oracle NKP supervision. This decouples high-level navigational intent from low-level motion dynamics, yielding a well-conditioned learning problem and preventing gradient interference when subsequently estimating $p(Z\mid X)$. To capture the sequential structure of vessel motion, the future trajectory distribution is factorized autoregressively as
$p(Y\mid X,Z)=\prod_{t=1}^{T}p(y_{t}\mid X,Z,y_{<t}),$ (6)
and optimized using maximum likelihood with teacher forcing.
While autoregressive long-horizon prediction is expressive, it is susceptible to error accumulation due to exposure bias, arising from the mismatch between training-time conditioning on ground-truth prefixes and inference-time conditioning on model predictions. To mitigate this issue, we adopt a step-by-step training scheme, which stabilizes learning and reduces cumulative error propagation. From a reinforcement learning perspective, this procedure is equivalent to behavior cloning, where the model learns to imitate expert actions, i.e., SOG and COG of the vessels, conditioned on historical states and NKP information. These two training schemes were conducted alternatively with 50 epochs and 10 epochs, respectively. The loss functions for these two approaches are:
$\mathcal{L}_{GPT2}=\frac{1}{T}\sum^{T}_{t=1}(\hat{\emph{vol}}(t)-\emph{vol}(t))^{2}$ (7)
$\mathcal{L}_{BC}=\frac{1}{T}\sum^{T}_{t=1}(\hat{\emph{coord}}(t)-\emph{coord}(t))^{2}$ (8)
vol refers to the SOG component over latitude and longitude, while coord refers to the latitude and longitude. The latter loss function can be interpreted as minimizing a one-step imitation loss with respect to the expert policy induced by AIS trajectories. This model can be interpreted as the Maritime Trajectory Large Model for downstream tasks related to trajectory understanding.
3.3.2 NKP Modeling (green)
In the second stage, we focus on estimating the NKP posterior $p(Z\mid X)$. The trajectory prediction backbone trained in Stage 1 is frozen, and only the NKP prediction feedforward module is optimized using the contrastive loss in Eq. (2). Freezing the backbone stabilizes representation learning and prevents interference between low-level motion dynamics and high-level semantic intent, ensuring that the learned NKP embeddings capture coarse navigational information rather than fine-grained kinematic details. With that, NKP can be derived as a byproduct.
3.3.3 Integrated Inference
[…]
3.4 Model Architecture and Physical Representation
To implement this factorization, we build our model upon a decoder-only Transformer (Vaswani et al., 2017), named MiniMind [^1] https://github.com/jingyaogong/minimind, due to its efficiency and strong temporal modeling capabilities. We introduce linear layers as input and output projections in the Transformer to process continuous maritime data, mapping latitude and longitude to continuous embeddings and SOG/COG to scaled action components. Specifically, we normalize the longitude and latitude to [1,1][-1,1] and convert COG and SOG into latitude/longitude velocity components using a scaling factor of $1/25$, which can be interpreted as the combined effect of crew decisions and environmental dynamics. However, to visualize the motion, the next latitude and longitude should be calculated, given the predicted SOG and COG.
Assumption 3.2 (Local Linear Motion).
Over a short time interval $T$, the vessel is assumed to move with constant speed over ground (SOG) $v$ and course over ground (COG) $\theta$.
Proposition 3.3 (SOG/COG Coordinate Update).
Under 3.2, given current coordinates $(lat_{0},lon_{0})$, the next position $(lat_{1},lon_{1})$ is given by
$lat_{1}$ $=lat_{0}+\frac{v\cos\theta}{R}T,$ (9)
$lon_{1}$ $=lon_{0}+\tan\theta\ln\frac{\lvert\sec(lat_{1})+\tan(lat_{1})\rvert}{\lvert\sec(lat_{0})+\tan(lat_{0})\rvert}$, (10)
where $R$ denotes the Earth’s radius.
The derivation follows from integrating the velocity field under a spherical-Earth approximation and is presented in Appendix B. Fig. 3 clearly indicates that our assumption holds, with purple and green points representing prediction and ground truth, respectively.
Figure 3: from COG & SOG to latitude & longitude - dense. Purple and green points represent the next coordinates derived by the formula and ground truth, respectively. The fact that the purple and green dots are close enough indicates the practical validity of the formulation.
Table 1: Qualitative comparison of MP-LSTM, TrAISformer, and our model. They are trained on the same training split with the same procedure, and evaluated on the same test split.
Ours [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MP-LSTM [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
trAISformer [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 2: Ablation Experiment for NKP Prediction. 6ch refers to the model with correct NKP, while 6ch w/ WD refers to the model with random NKP.
4ch [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
6ch [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
6ch w/ WD [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
4 Evaluation Metrics
Mean Squared Error of Position (MSEP).
We evaluate point-level trajectory accuracy using the Mean Squared Error of Position (MSEP), defined as
$\mathrm{MSEP}=\frac{1}{T}\sum_{t=1}^{T}\left\|P^{\mathrm{pred}}(t)-P^{\mathrm{true}}(t)\right\|_{2}^{2},$ (11)
where $T$ denotes the prediction horizon, and $P^{\mathrm{pred}}(t)$ and $P^{\mathrm{true}}(t)$ represent the predicted and ground-truth positions at time step $t$, respectively.
MSEP measures the point-wise positional discrepancy between predicted and ground-truth trajectories and thus reflects local prediction accuracy. A smaller MSEP indicates more accurate trajectory predictions at the coordinate level.
Mean Squared Error of Curvature (MSEC).
To evaluate trajectory smoothness rather than geometric curvature itself, we measure discrepancies in curvature profiles between predicted trajectories and the ground truth. We define the curvature at each trajectory point as
$\kappa_{i}=\begin{cases}0,&i\in\{0,n-1\},\\ \frac{\Delta\theta_{i}}{\bar{d}_{i}},&\text{otherwise},\end{cases}$ (12)
where $\Delta\theta_{i}$ denotes the change in heading angle at point $i$, and $\bar{d}_{i}$ is the average arc length of the adjacent segments. To reduce local noise, a smoothed curvature is computed using a three-point moving average:
$\kappa_{i}^{\mathrm{smooth}}=\frac{\kappa_{i-1}+\kappa_{i}+\kappa_{i+1}}{3}.$ (13)
Based on the smoothed curvature, we define the Mean Squared Error of Curvature (MSEC) between the predicted trajectory and the ground truth as
$\mathrm{MSEC}=\frac{1}{n}\sum_{i=0}^{n-1}\left(\kappa_{i}^{\mathrm{pred}}-\kappa_{i}^{\mathrm{true}}\right)^{2}.$ (14)
The MSEC evaluates trajectory quality at a microscopic level by measuring point-wise curvature discrepancies and thus reflects the smoothness of the predicted trajectory. For MP-LSTM, only two trajectory points are directly predicted and the remaining points are obtained via interpolation, resulting in a degenerate curvature profile with zero curvature variation. Consequently, MSEC does not apply to MP-LSTM in Table 3. The curvature definition follows the standard discrete approximation $\kappa\approx d\theta/ds$ commonly used in trajectory analysis and motion planning (LaValle, 2006).
Mean Fréchet Distance (MFD).
The Mean Fréchet Distance (MFD) (Alt and Godau, 1995) is defined as
$\mathrm{MFD}=\frac{1}{B}\sum_{k=1}^{B}F(A_{k},B_{k}),$ (15)
where $F(\cdot,\cdot)$ denotes the discrete Fréchet distance between a predicted trajectory and its ground truth. This metric evaluates curve-level performance.
Evaluation Metrics and Dataset Overview.
To comprehensively evaluate trajectory prediction performance at different granularities, we adopt three complementary metrics: Mean Squared Error of Position (MSEP), Mean Squared Error of Curvature (MSEC), Mean Fréchet Distance (MFD) and inference duration. Specifically, MSEP measures point-wise positional accuracy and reflects local coordinate-level errors. MSEC evaluates discrepancies in curvature profiles and characterizes the smoothness and local geometric consistency of predicted trajectories. In contrast, MFD captures curve-level similarity by assessing the global geometric alignment between predicted and ground-truth trajectories. The inference durations, in seconds, evaluate the model’s efficiency and include the time taken on the same test datasets at the same batch sizes. Together, these metrics provide a multi-scale evaluation framework that jointly accounts for local accuracy, trajectory smoothness, global shape consistency, and time efficiency. More datasets and implementation details were covered in Appendix Appendix C and Appendix D.
5 Results
5.1 Metrics on Trajectory Prediction
Table 3 reports a quantitative comparison between our method and two representative baselines, MP-LSTM [^2] We reapplied it and code are in supplementary materials. (Gao et al., 2021) and TrAISformer [^3] https://github.com/CIA-Oceanix/TrAISformer (Nguyen and Fablet, 2024) , evaluated using four complementary metrics: MSEP, MSEC, MFD, and inference time.
Overall, our method achieves the best performance across all reported metrics. Specifically, our model attains the lowest MSEP, indicating more accurate point-wise predictions over the entire time horizon. At the curve level, our approach also achieves the smallest MFD, reflecting closer geometric and curvature-level alignment with the ground-truth trajectories. In addition, our model demonstrates a clear advantage in inference efficiency. As shown in Table 3, it requires significantly less inference time than both MP-LSTM and TrAISformer, making it more suitable for large-scale and real-time vessel trajectory prediction scenarios.
Table 3: Quantitative comparison of trajectory prediction performance between our method, MP-LSTM, and TrAISformer, evaluated using MSEP, MSEC, MFD, and inference time on the whole test dataset.
metrics MSEP MSEC MFD time
our model 0.41 1.23e-03 7.80 62
MP-LSTM 1.6 0.00 31.11 287
TrAISformer 0.71 1.19e-02 19.78 541
5.2 Generalization to Public AIS Dataset
We further evaluate the generalization capability of our method on the public AIS dataset (Xie et al., 2025), which is collected from a different source and exhibits a distribution distinct from our private dataset.
None of the evaluated models are trained or fine-tuned on this dataset, and all results therefore reflect true out-of-domain generalization performance. As shown in Table 4, our method achieves the best performance across all reported metrics, maintaining both low prediction error and high inference efficiency. In contrast, TrAISformer exhibits severe performance degradation on this dataset, as its reliance on embedding layers prevents it from generalizing to unseen latitude-longitude trajectories.
Table 4: Metrics for trajectory prediction on the public AIS_Dataset (Xie et al., 2025).
metrics MSEP MSEC MFD time
our model 2.8e-04 2.33e-07 0.96 20
MP-LSTM 3.83e-04 0.00 2.77 52
TrAISformer 2.04 9.76e-03 49.50 121
5.3 Next Key Point Prediction Accuracy
This section evaluates the effectiveness of our Next Key Point (NKP) prediction models. We compare three variants: pretrain-c, the MiniMind model trained from scratch; sft-c, a closed-set probabilistic fine-tuned MLP; and sft-o-s2, an open-set supervised fine-tuned model capable of extending to unseen key points derived by contrastive learning.
As shown in Table 5, all our proposed models significantly outperform the Random Forest baseline (Zhang et al., 2020). Among them, pretrain-c achieves the highest closed-set accuracy (98.66%), while sft-c also performs strongly (95.44%). The open-set fine-tuned sft-o-s2 attains slightly lower accuracy (93.61%) but offers the advantage of seamless generalization to new NKPs through simple database extension, without retraining. This demonstrates both the effectiveness of the fine-tuning procedure and its potential for application to other evolving prediction tasks.
Considering the trade-off between accuracy and flexibility, we adopt sft-o-s2 for integrated inference in subsequent experiments, as it provides robust performance while supporting scalable, open-set NKP prediction.
Table 5: Next Key Point (NKP) prediction accuracy. Tested trajectories include NKPs absent from the sft-o-s2 database.
Method Accuracy
sft-o-s2 93.61%
sft-c 95.44%
pretrain-c 98.66%
Random Forest(Zhang et al., 2020) 76.54%
5.4 Effect of NKP Quality on Trajectory Prediction
Table 6: Ablation of NKP Prediction. Correct and wrong NKP are oracle and adversarial inputs, while SFT-O, SFT-C, and pretrain-C are Stage 2 NKP prediction strategies. NKPs that do not appeal in the database are omitted.
metrics MSEP MSEC MFD
correct NKP 0.41 9.48e-04 7.78
SFT-O 0.41 1.23e-03 7.80
SFT-c 0.41 5.58e-02 8.71
wrong NKP 0.41 0.32 11.04
pretrain-c 0.36 1.01e-03 11.72
pure 4ch 0.41 1.04 8.99
We analyze the impact of Next Key Point (NKP) quality on downstream trajectory prediction in Table LABEL:tag:_stage_3_compare_with_approaches. The correct and wrong NKP settings define oracle upper and adversarial lower bounds, within which all models using predicted NKPs fall, indicating that NKP prediction errors induce only limited and acceptable degradation.
Among all strategies, sft-o achieves performance closest to the oracle across MSEP, MSEC, and MFD, while exhibiting the most stable behavior; together with its open-set inference capability, this motivates its selection as the final NKP prediction scheme. Although sft-c attains comparable point-wise accuracy, it introduces substantially larger curvature errors, suggesting that local accuracy alone does not guarantee smooth or coherent trajectories.
The pretrain-c model, which adopts the original MiniMind backbone without explicitly decoupling NKP encoding from motion modeling, exhibits significantly larger MFD despite lower MSEP, indicating a tendency toward overfitting due to non-temporal NKP injection into temporal models and increased model capacity.
Importantly, most NKP-conditioned models outperform the pure 4-channel baseline, confirming that incorporating NKPs as high-level intent provides critical global guidance beyond local motion cues, while incorrect NKPs cause severe trajectory distortion. Overall, these results demonstrate that NKP quality is crucial for achieving smooth and globally consistent long-horizon trajectory prediction.
5.5 Qualitative Results and Visualization
We qualitatively evaluate trajectory predictions against ground truth across different regions (Table 1). MP-LSTM (Gao et al., 2021), reimplemented using a training-set database, predicts destinations accurately but exhibits large deviations at support points. TrAISformer (Nguyen and Fablet, 2024), applied from the official repository, shows pauses and stagnant latitude. In contrast, our model accurately captures both straight and turning trajectories, maintaining correct timing and coordinates, which demonstrates robust long-horizon prediction in complex scenarios.
Table 2 highlights the effectiveness of the proposed NKP channel. Models with NKP consistently produce trajectories closer to the ground truth compared to those without, using the same random seed for fair comparison. The 6-channel row with an incorrect NKP input shows that the model effectively decouples NKP information: errors in NKP cause the predicted trajectory to twist, confirming the model’s sensitivity to NKP guidance while retaining robustness.
6 Conclusion
We presented a hierarchical framework for long-horizon vessel trajectory prediction that incorporates Next Key Points (NKPs) as explicit semantic intent. By decoupling high-level intent inference from conditional motion modeling, the framework reduces long-term uncertainty and improves trajectory coherence.
Beyond maritime navigation, this intent-conditioned factorization provides a general modeling paradigm for other long-horizon sequential prediction problems with hierarchical decision structures. Moreover, our approach can serve as a foundation for developing maritime AI solutions that can enable more effective trajectory understanding and offer a base representation for downstream tasks such as anomaly detection, route optimization, or multi-agent coordination.
Appendix A On Information Gain by Conditioning
In this appendix, we formally clarify the relationship between conditioning on additional information and the reduction of uncertainty. While conditioning on an auxiliary variable does not necessarily increase posterior probabilities pointwise, it always improves certainty in an information-theoretic or decision-theoretic sense.
A.1 Preliminaries
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and let $X,Y,Z$ be random variables defined on it. Denote by $\sigma(X)$ and $\sigma(X,Y)$ the $\sigma$-algebras generated by $X$ and $(X,Y)$ respectively, with $\sigma(X)\subseteq\sigma(X,Y)$.
We use $H(\cdot)$ to denote Shannon entropy and $I(\cdot;\cdot\mid\cdot)$ to denote conditional mutual information.
A.2 Conditioning and Uncertainty Reduction
We begin with a fundamental result stating that conditioning on additional information cannot increase uncertainty.
Lemma A.1 (Monotonicity of Conditional Entropy).
For any random variables $X,Y,Z$, the following inequality holds:
$H(Z\mid X,Y)\;\leq\;H(Z\mid X).$ (16)
Proof.
By the chain rule of entropy, we have
$H(Z\mid X)-H(Z\mid X,Y)=I(Z;Y\mid X),$ (17)
where $I(Z;Y\mid X)$ denotes the conditional mutual information. Since mutual information is non-negative, i.e., $I(Z;Y\mid X)\geq 0$, the result follows. ∎
Lemma A.1 provides a precise mathematical interpretation of the statement that conditioning on additional information increases certainty: the posterior uncertainty of $Z$ is reduced in an information-theoretic sense.
A.3 Why Pointwise Posterior Improvement Is Not Guaranteed
Despite the reduction in entropy, conditioning on $Y$ does not generally imply a pointwise increase of posterior probabilities.
Lemma A.2 (Tower Property of Conditional Expectation).
For any random variables $X,Y,Z$, we have
$\mathbb{E}\!\left[\,\mathbb{P}(Z\mid X,Y)\mid X\,\right]=\mathbb{P}(Z\mid X).$ (18)
Proof.
This follows directly from the tower property of conditional expectation:
$\mathbb{E}\!\left[\mathbb{E}[\mathbf{1}_{Z}\mid X,Y]\mid X\right]=\mathbb{E}[\mathbf{1}_{Z}\mid X]=\mathbb{P}(Z\mid X).$ (19)
Lemma A.2 implies that any pointwise monotonic inequality such as $\mathbb{P}(Z\mid X,Y)\geq\mathbb{P}(Z\mid X)$ cannot hold almost surely unless equality holds everywhere. Therefore, the benefit of conditioning should not be interpreted as a pointwise increase in posterior probabilities.
A.4 Decision-Theoretic Interpretation
The information gain induced by conditioning admits an equivalent interpretation in decision theory.
Theorem A.3 (Bayes Risk Monotonicity).
Let $L$ be any loss function. Then
[…]
Proof.
Any decision rule based on $X$ alone is a special case of a decision rule based on $(X,Y)$ that ignores $Y$. Taking the infimum over a larger class of decision rules cannot increase the minimum achievable risk. ∎
Theorem A.3 formalizes the intuition that additional information cannot worsen optimal decision performance.
A.5 Remarks on Stronger Monotonicity Notions
We emphasize that stronger notions of monotonicity, such as pointwise improvement of posterior probabilities or monotonicity in likelihood ratio order, require additional structural assumptions on the conditional distribution $P(Y\mid Z,X)$. Without such assumptions, uncertainty reduction should be understood exclusively in an information-theoretic or decision-theoretic sense, as established above.
Appendix B Detailed Derivation of Next-Timestamp Coordinates
Given current coordinates $lat_{0},lon_{0}$, current SOG $v$, and COG $\theta$, with a given short time period $T$, the new vessel coordinates $lat_{1},lon_{1}$ were expected to be calculated. Assume that the vessels perform uniform linear motion in the short time period. Then the coordinates in the next timestamp are:
Then, in the latitudinal direction, the SOG components are:
$v_{lat}=v\ \cos\theta$ (21)
As the radii remain,
$\omega_{lat}=\frac{v_{lat}}{R}$ (22)
And the changes in longitude are,
$\Delta lat=\omega_{lat}t=\frac{vt\cos\theta}{R}$ (23)
The relationships between latitude and time are
$lat(t)=lat_{0}+\frac{v\cos\theta}{R}t$ (24)
In the longitudinal direction, the SOG components are:
$v_{lat}=v\sin\theta$ (25)
The relationships between radii and latitude are
$r=R\cos(lat)$ (26)
The angular velocity is
$\omega_{lon}=\frac{v\sin\theta}{R\cos(lat)}$ (27)
Take (24) into (27) that
$\omega_{lot}(t)=\frac{v\sin\theta}{R\cos(lat)}=\frac{v\sin\theta}{R\cos(lat_{0}+\frac{v\cos\theta}{R}t)}$ (28)
Then the changes of longitudes are
$\Delta lon=\int^{t}_{0}\omega_{lon}(t)dt=\frac{v\sin\theta}{R}\int^{t}_{0}\frac{dt}{\cos(lat_{0}+\frac{v\cos\theta}{R}t)}$ (29)
With the formula
$\int\frac{1}{\cos\theta}d\theta=\ln|\sec\theta+\tan\theta|+C$ (30)
We can derive
$\Delta lon=\frac{Rv\sin\theta}{v\cos\theta R}\int^{t}_{0}\frac{1}{\cos(lat_{0}+\frac{v\cos\theta}{R}t)}d(\frac{v\cos\theta}{R})$ (31)
$=\tan\theta(\ln|\sec\theta+\tan\theta|)|^{\frac{v\cos\theta}{R}+lat_{0}}_{lat_{0}}$ (32)
Denoting $lat_{1}=lat_{0}+\frac{v\cos\theta}{R}T$
$\Delta lon=\tan\theta\ln\frac{|\sec(lat_{1})+\tan(lat_{1})|}{|\sec(lat_{0})+\tan(lat_{0})|}$ (33)
Finally, we can conclude that the new coordinates are
$lat_{1}=lat_{0}+\frac{v\cos\theta}{R}T$ (34)
$lon_{1}=lon_{0}+\tan\theta\ln\frac{|\sec(lat_{1})+\tan(lat_{1})|}{|\sec(lat_{0})+\tan(lat_{0})|}$ (35)
Appendix C Dataset and Data Preprocessing
C.1 Data Description and Coverage for Private Dataset
The dataset is proprietary and cannot be publicly released due to licensing restrictions. However, all baseline methods are trained and evaluated on the same dataset under identical preprocessing and evaluation protocols. Each data point includes the current timestamp, current longitude and latitude, current Speed Over Ground (SOG), Course Over Ground (COG), and the destination port name with its corresponding coordinates.
Our dataset contains comprehensive global maritime coverage with the following geographical boundaries:
  • Latitude range: $36.4^{\circ}S$ to $55.5^{\circ}N$
  • Longitude range: $180.0^{\circ}W$ to $180.0^{\circ}E$
This extensive spatial coverage includes major shipping routes across the Atlantic, Pacific, and Indian Oceans, as well as significant regional seas, ensuring the model’s exposure to diverse maritime navigation patterns. All vessels in this private dataset are ocean-going container ships engaged in long-haul international trade.
C.2 Data Preprocessing Pipeline
The data preprocessing consists of two steps:
Step 1 - Temporal Interpolation:
Due to the non-uniform temporal distribution with AIS system limitation, we perform interpolation with a fixed time interval $T=5$ minutes to achieve uniform sampling.
Step 2 - Spatial Annotation:
The Next Key Point is determined by intersecting each trajectory with predefined maritime key nodes (e.g., ports and straits), and all points between two consecutive intersections are labeled with the corresponding NKP, yielding 103 unique spatial annotations.
After uniform interpolation, the dataset comprises 15,728,640 waypoints and corresponds to 1,310,720 hours of maritime navigation data, providing a substantial foundation for model training. The model received the data with windows sliding within different MMSI codes. The train, validation, and test datasets were split by MMSI code and are free of data leakage.
C.3 Stage 2 Database Sampling Strategy
To sample the database for open-set verification, the dataset is partitioned into multiple voyage segments based on Maritime Mobile Service Identity (MMSI) and Key Points. Within each segment, we apply a sliding-window approach to extract trajectory sequences, with a window size of $L_{seq}$ and a stride of $S$ steps, where $L_{seq}$ denotes the sequence length for model input. In practice, we set it as 288, i.e., 48 hours.
After the sliding-window operation, we perform stratified sampling by randomly selecting 50 trajectories per key node, achieving a good balance between representation diversity and computational efficiency. Key Points with insufficient samples (fewer than 50 points) are excluded from the final dataset. This methodology yields a balanced dataset comprising 56 ports and 2,800 data entries, ensuring adequate representation across different geographical regions.
C.4 Data Description and Coverage for Public Dataset
We further evaluate our method on a public AIS dataset published on Hugging Face to assess generalization beyond our private data source. Each AIS record contains the vessel identifier (MMSI), timestamp (BaseDateTime), geolocation (latitude and longitude), kinematic attributes (speed over ground, course over ground, and heading), vessel type encoded using standard AIS ship-type codes, navigation status, and a track identifier when available.
The version used in our experiments exhibits broad spatial and kinematic coverage.
  • Latitude range: $0.05^{\circ}\ N$ to $61.25^{\circ}\ N$
  • Longitude range: $177.86^{\circ}\ W$ to $171.22^{\circ}\ E$
  • SOG has a mean of $11.98\pm 4.39$ knots with the range $[1.00,29.90]$
  • COG has a mean of $185.73\pm 101.72$ with the range cover the full angular domain $0^{\circ},360^{\circ})$
Overall, the dataset predominantly covers near-equatorial to mid- and high-latitude regions in the Northern Hemisphere and includes trajectories across major ocean basins. It is qualified as a test dataset.
The public dataset contains multiple vessel categories, including Cargo, Fishing, Passenger, Pleasure Craft/Sailing, Tanker, and Tug/Tow. To ensure consistency with our private dataset and experimental focus, we restrict evaluation to cargo vessels by selecting AIS messages whose VesselType falls within the cargo ship-type codes (70–79).
According to the repository statistics, the dataset comprises AIS messages from approximately 19k vessels with hundreds of millions of records. In the snapshot used for evaluation, we observe 19,016 unique MMSIs, with individual vessel trajectories spanning from several days to multiple months, supporting both short- and long-horizon prediction scenarios.
To align preprocessing with our private dataset, we apply a unified pipeline: AIS messages are filtered by vessel type, grouped by MMSI (and TrackID when available), and resampled via linear interpolation to a fixed temporal resolution of $T=5$ minutes. All features are cast to consistent units and data types to ensure compatibility across datasets. The process details are the same as the private dataset. The private test dataset has 2513 samples, while the public test dataset has 560 samples and comes from distinct sources.
Appendix D Implementation Details
D.1 Model Implementation
The hidden size of MiniMind Block was set as 256 with 2 key-value heads. MiniMind Model 1 has 3 hidden layers, while MiniModel 2 has 1. The threshold of cosine similarity is set as 0.5, and the hidden size of the MLP is 128.
D.2 Training Details
The learning rate is 7e-5. We also utilized AdamW as the optimizer and Cosine Annealing Warm Restarts as the scheduler. Other hyperparameters are default. All training and evaluation are performed on a single NVIDIA RTX 4090.
A portion of the code, model checkpoints, and a processed public dataset have been attached to the supplemental materials. The whole code and checkpoints will be public after acceptance.
Appendix E Broader Impact
This work studies long-horizon trajectory prediction for maritime vessels based on historical navigation data. Potential positive impacts include supporting strategic-level maritime transportation analysis, such as providing voyage-status estimates for ship operators and cargo owners, and enabling long-term route and schedule assessment under varying environmental conditions. These applications may improve operational planning and overall efficiency in maritime logistics.
The proposed method focuses on offline prediction over extended time horizons and is not designed for real-time tracking or operational decision-making. The data used in this study consist of historical vessel trajectories and do not involve personal or sensitive information.

### v2

@@
 Nevertheless, we observe that standard autoregressive Transformers still degrade severely in long-horizon settings.
-Notably, both TrAISformer and recent token-based GPT-style forecasters, such as TrackGPT (Stroh, 2024), convert continuous trajectory attributes into discrete tokens and learn embeddings over these discrete representations as the input to a Transformer backbone, which will prevent unseen embedded layer regions from being trained and will be fail when trajectories enter those regions.
+Notably, both TrAISformer (Nguyen and Fablet, 2024) and recent token-based GPT-style forecasters, such as TrackGPT (Stroh, 2024), convert continuous trajectory attributes into discrete tokens and learn embeddings over these discrete representations as the input to a Transformer backbone.
+Due to the discretization of continuous geographic space, token-based trajectory models are trained only on a sparse subset of spatial tokens.
+When deployed in previously unseen regions, the model must extrapolate beyond the empirical support of the training distribution, often leading to unstable or incoherent predictions.
 As uncertainty accumulates, predictions tend to collapse toward nearly horizontal trajectories.
@@
 The loss function is:
-$\mathcal{L}_{\text{TCL}}=\frac{1}{B}\sum_{i=1}^{B}\big[(1-y_{i})D_{i}^{2}+y_{i}\max(0,M-D_{i})^{2}\big],$ (3)
-where $y_{i}=1$ if the pair shares the same NKP label, otherwise $y_{i}=0$, and $M$ is a margin.
 Figure 1: Overall Architecture.
@@
 Then, hidden states 1 and 2 will be concatenated and passed through a dense layer, MiniMind model 2, and a decoder to obtain the predicted SOG and COG.
+$\mathcal{L}_{\text{TCL}}=\frac{1}{B}\sum_{i=1}^{B}\big[(1-y_{i})D_{i}^{2}+y_{i}\max(0,M-D_{i})^{2}\big],$ (3)
+where $y_{i}=1$ if the pair shares the same NKP label, otherwise $y_{i}=0$, and $M$ is a margin.
 3 Methodology
@@
 3.2 Semantic Next-Key-Point Modeling
-Definition 3.1 (Semantic Navigational Key Point (NKP)).
-A Semantic Navigational Key Point (NKP) is defined as a latent variable that represents high-level navigational intent of a vessel, rather than a deterministic geometric waypoint.
-Specifically, an NKP corresponds to an equivalence class of future trajectories that share the same navigational semantics, such as passing through a specific channel, approaching a port region, or following a major shipping lane.
+Definition 3.1 (Semantic Next Key Point (NKP)).
+A Semantic Next Key Point (NKP) is a latent semantic variable that represents the anticipated next navigational anchor of a vessel, inferred from its observed trajectory history.
+Unlike a deterministic geometric waypoint, an NKP encodes high-level future intent by characterizing an equivalence class of plausible future trajectories that share consistent navigational semantics, such as heading toward a port region, passing through a specific strait, or following a major shipping lane.
+NKPs are not the terminal points, and by introducing NKPs, an extremely long trajectory can be decomposed into several shorter segments, each governed by a coherent semantic intent, enabling hierarchical long-horizon prediction.
+Remark.
+We use the term semantic to emphasize that an NKP represents an intent-level equivalence class of feasible futures rather than a precise geometric waypoint, although in practice we obtain NKP supervision from spatial annotations (e.g., port/strait intersections) for training and evaluation.
 To incorporate Next Key Point (NKP) information, we explicitly estimate the conditional distribution $p(Z\mid X)$, where $X$ denotes the observed trajectory history and $Z$ the corresponding NKP.
@@
 In practice, we approximate marginalization over $Z$ using a MAP estimate, which performs robustly in our experiments.
+When $P(Z\mid X)$ is sharply peaked, marginalization can be well approximated by conditioning on the MAP NKP.
 During inference, an unknown trajectory is compared against a database of reference trajectories.
+The reference database contains only training trajectories for both the public and private test datasets.
 For each reference trajectory whose similarity exceeds a predefined threshold, a vote is cast for its associated NKP.
 The NKP with the highest vote count is selected as the prediction, as summarized in Algorithm 1.
-This voting procedure can be interpreted as a non-parametric approximation of $p(Z\mid X)$ through nearest-neighbor evidence aggregation.
+This voting procedure can be interpreted as a nonparametric approximation to $p(Z\mid X)$ via nearest-neighbor evidence aggregation.
 The overall training and inference pipeline is shown in Fig.
@@
 where $R$ denotes the Earth’s radius.
-The derivation follows from integrating the velocity field under a spherical-Earth approximation and is presented in Appendix B.
-Fig.
-3 clearly indicates that our assumption holds, with purple and green points representing prediction and ground truth, respectively.
-Figure 3: from COG & SOG to latitude & longitude - dense.
-Purple and green points represent the next coordinates derived by the formula and ground truth, respectively.
-The fact that the purple and green dots are close enough indicates the practical validity of the formulation.
-Table 1: Qualitative comparison of MP-LSTM, TrAISformer, and our model. They are trained on the same training split with the same procedure, and evaluated on the same test split.
-Ours [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-MP-LSTM [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-trAISformer [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Table 2: Ablation Experiment for NKP Prediction. 6ch refers to the model with correct NKP, while 6ch w/ WD refers to the model with random NKP.
-4ch [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-6ch [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-6ch w/ WD [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+This formulation operates in a locally Euclidean space and avoids repeated spherical projections, thereby improving numerical stability and preventing systematic error amplification during long-horizon autoregressive rollout.
+The derivation follows from integrating the velocity field under a spherical-Earth approximation and is presented in Appendix C, along with a numerical stability analysis in Appendix D.
+To empirically assess the numerical consistency of the proposed coordinate update, we perform a one-step displacement check on real AIS trajectories.
+For each trajectory segment of length $288$, the next-step position is computed from the current SOG/COG and compared against the ground-truth next coordinate.
+Across the evaluated trajectories, the resulting mean squared error is on the order of $10^{-9}$, indicating numerical consistency up to machine precision.
 4 Evaluation Metrics
@@
 Together, these metrics provide a multi-scale evaluation framework that jointly accounts for local accuracy, trajectory smoothness, global shape consistency, and time efficiency.
-More datasets and implementation details were covered in Appendix Appendix C and Appendix D.
+More datasets and implementation details were covered in Appendix Appendix E and Appendix F.
+Table 1: Qualitative comparison of MP-LSTM, TrAISformer, and our model.
+They are trained on the same training split with the same procedure, and evaluated on the same test split.
+Blue, green, red, orange, and purple trajectories refer to input, ground truth, our prediction, prediction from MP-LSTM (Gao et al., 2021), and prediction from TrAISformer (Nguyen and Fablet, 2024), respectively.
+Additional qualitative examples are provided in Appendix B.
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 5 Results
@@
 Table 2: Quantitative comparison of trajectory prediction performance between our method, MP-LSTM, and TrAISformer, evaluated using MSEP, MSEC, MFD, and inference time on the whole test dataset.
+The unit of time is second, and was tested on an NVIDIA RTX 4090 for the whole test dataset with 2513 samples, and a batch size of 128.
 metrics MSEP MSEC MFD time
@@
 We further evaluate the generalization capability of our method on the public AIS dataset (Xie et al., 2025), which is collected from a different source and exhibits a distribution distinct from our private dataset.
-None of the evaluated models are trained or fine-tuned on this dataset, and all results therefore reflect true out-of-domain generalization performance.
-As shown in Table 4, our method achieves the best performance across all reported metrics, maintaining both low prediction error and high inference efficiency.
-In contrast, TrAISformer exhibits severe performance degradation on this dataset, as its reliance on embedding layers prevents it from generalizing to unseen latitude-longitude trajectories.
-Table 4: Metrics for trajectory prediction on the public AIS_Dataset (Xie et al., 2025).
+Table 3: Metrics for trajectory prediction on the public AIS_Dataset (Xie et al., 2025). All the metrics are the same as Table 2 with a total of 560 samples.
 metrics MSEP MSEC MFD time
@@
 TrAISformer 2.04 9.76e-03 49.50 121
+None of the evaluated models are trained or fine-tuned on this dataset, and all results therefore reflect true out-of-domain generalization performance.
+As shown in Table 3, our method achieves the best performance across all reported metrics, maintaining both low prediction error and high inference efficiency.
+In contrast, TrAISformer exhibits severe performance degradation on this dataset, as its reliance on embedding layers prevents it from generalizing to unseen latitude-longitude trajectories.
 5.3 Next Key Point Prediction Accuracy
@@
 We compare three variants: pretrain-c, the MiniMind model trained from scratch; sft-c, a closed-set probabilistic fine-tuned MLP; and sft-o-s2, an open-set supervised fine-tuned model capable of extending to unseen key points derived by contrastive learning.
-As shown in Table 5, all our proposed models significantly outperform the Random Forest baseline (Zhang et al., 2020).
-Among them, pretrain-c achieves the highest closed-set accuracy (98.66%), while sft-c also performs strongly (95.44%).
-The open-set fine-tuned sft-o-s2 attains slightly lower accuracy (93.61%) but offers the advantage of seamless generalization to new NKPs through simple database extension, without retraining.
-This demonstrates both the effectiveness of the fine-tuning procedure and its potential for application to other evolving prediction tasks.
-Considering the trade-off between accuracy and flexibility, we adopt sft-o-s2 for integrated inference in subsequent experiments, as it provides robust performance while supporting scalable, open-set NKP prediction.
 Table 4: Next Key Point (NKP) prediction accuracy. Tested trajectories include NKPs absent from the sft-o-s2 database.
@@
 sft-c 95.44%
-pretrain-c 98.66%
+pretrain-c 98.98%
 Random Forest(Zhang et al., 2020) 76.54%
+As shown in Table 4, all our proposed models significantly outperform the Random Forest baseline (Zhang et al., 2020).
+Among them, pretrain-c achieves the highest closed-set accuracy (98.98%), while sft-c also performs strongly (95.44%).
+The open-set fine-tuned sft-o-s2 attains slightly lower accuracy (93.61%) but offers the advantage of seamless generalization to new NKPs through simple database extension, without retraining.
+This demonstrates both the effectiveness of the fine-tuning procedure and its potential for application to other evolving prediction tasks.
+Considering the trade-off between accuracy and flexibility, we adopt sft-o-s2 for integrated inference in subsequent experiments, as it provides robust performance while supporting scalable, open-set NKP prediction.
 5.4 Effect of NKP Quality on Trajectory Prediction
+We analyze the impact of Next Key Point (NKP) quality on the prediction of downstream trajectory in Table 5.
+The correct and wrong NKP settings define oracle upper and adversarial lower bounds, within which all models using predicted NKPs fall, indicating that NKP prediction errors induce only limited and acceptable degradation.
 Table 5: Ablation of NKP Prediction.
@@
 pure 4ch 0.41 1.04 8.99
-We analyze the impact of Next Key Point (NKP) quality on downstream trajectory prediction in Table LABEL:tag:_stage_3_compare_with_approaches.
-The correct and wrong NKP settings define oracle upper and adversarial lower bounds, within which all models using predicted NKPs fall, indicating that NKP prediction errors induce only limited and acceptable degradation.
+Table 6: Ablation Experiment for NKP Prediction.
+4ch refers to the model without NKP information, 6ch refers to the model with correct NKP, while 6ch w/ WD refers to the model with random but wrong NKP.
+4ch [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+6ch [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+6ch w/ WD [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 Among all strategies, sft-o achieves performance closest to the oracle across MSEP, MSEC, and MFD, while exhibiting the most stable behavior; together with its open-set inference capability, this motivates its selection as the final NKP prediction scheme.
@@
 Importantly, most NKP-conditioned models outperform the pure 4-channel baseline, confirming that incorporating NKPs as high-level intent provides critical global guidance beyond local motion cues, while incorrect NKPs cause severe trajectory distortion.
-Overall, these results demonstrate that NKP quality is crucial for achieving smooth and globally consistent long-horizon trajectory prediction.
+Overall, these results demonstrate that the quality of NKP is crucial to achieving smooth and globally consistent long-horizon trajectory prediction.
 5.5 Qualitative Results and Visualization
-We qualitatively evaluate trajectory predictions against ground truth across different regions (Table 1).
-MP-LSTM (Gao et al., 2021), reimplemented using a training-set database, predicts destinations accurately but exhibits large deviations at support points.
+We qualitatively evaluate trajectory predictions against the ground truth across different regions (Table 1).
+MP-LSTM (Gao et al., 2021) predicts trajectories using only two anchor points (support and destination).
+While the support point is generally more accurate, the destination is often less reliable, and an error in either anchor can induce a significant global deviation in the reconstructed trajectory.
 TrAISformer (Nguyen and Fablet, 2024), applied from the official repository, shows pauses and stagnant latitude.
@@
 Table 6 highlights the effectiveness of the proposed NKP channel.
-Models with NKP consistently produce trajectories closer to the ground truth compared to those without, using the same random seed for fair comparison.
+Models with NKP consistently produce trajectories closer to the ground truth compared to those without it, using the same random seed for fair comparison.
 The 6-channel row with an incorrect NKP input shows that the model effectively decouples NKP information: errors in NKP cause the predicted trajectory to twist, confirming the model’s sensitivity to NKP guidance while retaining robustness.
@@
 Without such assumptions, uncertainty reduction should be understood exclusively in an information-theoretic or decision-theoretic sense, as established above.
-Appendix B Detailed Derivation of Next-Timestamp Coordinates
+Appendix B Additional Qualitative Results
+In this appendix, we provide additional qualitative comparisons between MP-LSTM, TrAISformer, and our method on diverse navigation scenarios.
+Table 7: Additional qualitative trajectory prediction examples.
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Appendix C Detailed Derivation of Next-Timestamp Coordinates
 Given current coordinates $lat_{0},lon_{0}$, current SOG $v$, and COG $\theta$, with a given short time period $T$, the new vessel coordinates $lat_{1},lon_{1}$ were expected to be calculated.
@@
 $lon_{1}=lon_{0}+\tan\theta\ln\frac{|\sec(lat_{1})+\tan(lat_{1})|}{|\sec(lat_{0})+\tan(lat_{0})|}$ (35)
-Appendix C Dataset and Data Preprocessing
-C.1 Data Description and Coverage for Private Dataset
+Appendix D Numerical Stability of Autoregressive Rollout
+We analyze the numerical behavior of autoregressive trajectory rollout under different coordinate update formulations.
+Let $s_{t}\in\mathbb{R}^{d}$ denote the true vessel state at time $t$ and $\hat{s}_{t}$ the predicted state, with prediction error $e_{t}=\hat{s}_{t}-s_{t}$.
+In autoregressive forecasting, predictions are propagated according to
+$\hat{s}_{t+1}=f(\hat{s}_{t}),$ (36)
+which induces the first-order error dynamics
+$e_{t+1}\approx J_{f}(s_{t})\,e_{t},$ (37)
+where $J_{f}$ denotes the Jacobian of the state transition function evaluated at $s_{t}$. Long-horizon numerical stability therefore depends critically on the spectral properties of $J_{f}$.
+D.1 Spherical Geometry–Based Updates
+In spherical or geodesic formulations, state updates are typically implemented as a composition of coordinate projection and motion integration:
+$f_{\mathrm{sph}}=\Pi^{-1}\circ G\circ\Pi,$ (38)
+where $\Pi$ maps geographic coordinates to a local tangent plane, and $G$ performs motion updates in that plane.
+The Jacobians of $\Pi$ and $\Pi^{-1}$ depend nonlinearly on latitude and heading, introducing location-dependent scaling factors. The resulting error propagation becomes
+$e_{t+1}\approx J_{\Pi^{-1}}\,J_{G}\,J_{\Pi}\,e_{t}.$ (39)
+Even if $J_{G}$ is well-conditioned, repeated application of $J_{\Pi}$ and $J_{\Pi^{-1}}$ across time steps leads to multiplicative error accumulation.
+As a result, small numerical inaccuracies can be systematically amplified during long-horizon rollout, especially at high latitudes or under frequent heading changes.
+D.2 Locally Euclidean Update (Ours)
+In contrast, our formulation performs state updates directly in a locally Euclidean coordinate system consistent with AIS-reported SOG and COG.
+Under the local linear motion assumption, the update function is smooth and approximately affine over short time intervals:
+$f_{\mathrm{ours}}(s_{t})=s_{t}+\Delta t\cdot v(s_{t}),$ (40)
+where $v(\cdot)$ denotes the velocity field induced by SOG and COG.
+The corresponding Jacobian satisfies
+$J_{f}\approx I+\mathcal{O}(\Delta t),$ (41)
+where $I$ is the identity matrix. Consequently, prediction errors accumulate approximately additively rather than multiplicatively:
+$\|e_{T}\|\leq\|e_{0}\|+\mathcal{O}(T\Delta t).$ (42)
+This behavior contrasts sharply with spherical formulations, where the effective Jacobian deviates from identity due to repeated nonlinear projections.
+Therefore, our locally Euclidean update is numerically more stable under long-horizon autoregressive rollout and prevents systematic drift caused by geometric distortion.
+Appendix E Dataset and Data Preprocessing
+E.1 Data Description and Coverage for Private Dataset
 The dataset is proprietary and cannot be publicly released due to licensing restrictions.
@@
 All vessels in this private dataset are ocean-going container ships engaged in long-haul international trade.
-C.2 Data Preprocessing Pipeline
+E.2 Data Preprocessing Pipeline
 The data preprocessing consists of two steps:
@@
 The train, validation, and test datasets were split by MMSI code and are free of data leakage.
-C.3 Stage 2 Database Sampling Strategy
+E.3 Stage 2 Database Sampling Strategy
 To sample the database for open-set verification, the dataset is partitioned into multiple voyage segments based on Maritime Mobile Service Identity (MMSI) and Key Points.
@@
 This methodology yields a balanced dataset comprising 56 ports and 2,800 data entries, ensuring adequate representation across different geographical regions.
-C.4 Data Description and Coverage for Public Dataset
+E.4 Data Description and Coverage for Public Dataset
 We further evaluate our method on a public AIS dataset published on Hugging Face to assess generalization beyond our private data source.
@@
 The private test dataset has 2513 samples, while the public test dataset has 560 samples and comes from distinct sources.
-Appendix D Implementation Details
-D.1 Model Implementation
+Appendix F Implementation Details
+F.1 Model Implementation
 The hidden size of MiniMind Block was set as 256 with 2 key-value heads.
@@
 The threshold of cosine similarity is set as 0.5, and the hidden size of the MLP is 128.
-D.2 Training Details
+F.2 Training Details
 The learning rate is 7e-5.
@@
 A portion of the code, model checkpoints, and a processed public dataset have been attached to the supplemental materials. The whole code and checkpoints will be public after acceptance.
-Appendix E Broader Impact
+Appendix G Broader Impact
 This work studies long-horizon trajectory prediction for maritime vessels based on historical navigation data.